Please help!!!, i need details explaination so i can write a lab report thanks a lottttttttttttttt. i'm working on a gravimatric sulfate experiment.

1. Does the solubility of BaSO4 increase significantly as the temperature of the solution is increase?. Look up the Ksp at different temperatures and calculate the solubility.

2. What are some errors in the procedure that could afft the final amount of precipitate?

3. From your answer to question 6, would you say that the procedure for the sulfate analysis is likely give results that are too high or too low?
Answer #1

1) No, the solubility is zero as it precipitates directly with increase in temperature generally we can say ppt. propotional to temp.

Note:

Ksp = [Ba2+] [CO32-j (s)(s) = s2 = (8.94 x 10-5)2= 8.0 x 10-9 =

2) Allowing the sample to boil than needed.

    While Barium sulphate to sample must take atleast 3 minutes for better result if more or less of mentioned time results in error.

  Ashless filter paper must be used other than this will give error in mass.

Don't let flames to ignite ashless filter paper while heating crucible instead it should combust on crucible heat unless it reduces mass of product.

3) By considering all the above errors error the sulphate analysis will give too low value in result.